OpenStudy (anonymous):

For the equation y = -3x2, find: a) Domain b) Range

OpenStudy (anonymous):

do u have to find the domain and range or is it asking if it is the domain or range?

OpenStudy (anonymous):

find the domain and range of: A)y=3x-9 Domain and Range are "All Real Numbers". --------------------------------------- B)y=-|x|+2 Domain: All Real Numbers Range: All Real Numbers <= 2 =============== C)y=2x^2+6x-7 Find the vertex: x = -b/(2a) = -6/(-4) = 3/4 Then f(3/4) = -30.38 ---- Domain: All Real Numbers Range:: All Real Numbers >= -4.061

OpenStudy (anonymous):

well nvr mind then

OpenStudy (anonymous):

sorry

OpenStudy (anonymous):

Can I get more help please?

OpenStudy (anonymous):

I can't get on that site. Wont let me.

OpenStudy (anonymous):

Usually the domain is all real numbers unless the function is a divide by zero. Your equation doesn't have any dividing, so your domain will be all real numbers

OpenStudy (anonymous):

I'm so lost..

OpenStudy (anonymous):

For the domain, that's the set of numbers that y can be. Looking at the function, you have a -3 and an x^2 That means it's a parabola which opens down with the vertex at the origin

OpenStudy (anonymous):

So the domain is >-3

OpenStudy (anonymous):

Can you ask me a question of what you don't understand?

OpenStudy (anonymous):

I don't understand the domain thing... or finding it.

OpenStudy (anonymous):

Ok! No problem! So the domain means all the x values you could plug in. So you could plug in 2 or 100 or -1 or 0 and you would still get a y value. But look at this type of equation Y=1/x If we plugged in 0 1/0 is not allowed in math, can't divide by zero. So your equation doesn't have any division, so that means all real numbers can be plugged in for x, or in other words, all real numbers for the domain.

OpenStudy (anonymous):

wait.. so whats range?

OpenStudy (anonymous):

The range is all the values that come out of the equation. Or all the possible values of y